OpenStudy (anonymous):

write a function rule to describe the statement, the amount of change c(x) from a $50 bill if you buy x pounds of chicken at $2.09 per pound

OpenStudy (anonymous):

c(x) = $50 - ($2.09) x how much change from 10 lb of chicken?

OpenStudy (anonymous):

$15.90?

OpenStudy (anonymous):

I got c(10) = 50 - 10(2.09) = 50 - 20.90 = 29.10 If you get that, you've probably understood the equation.

OpenStudy (anonymous):

will try again,ty

OpenStudy (anonymous):

Best wishes. You need to calculate the total cost of the chicken, then subtract that from the $50 you have.